# Restoring Calibration for Aligned Large Language Models: A Calibration-Aware Fine-Tuning Approach

### v1

@@ 1 Introduction
In this paper, we conduct a systematic investigation into two fundamental questions: (1) Why does preference alignment affect calibration? and (2) How can we effectively restore calibration while maintaining the benefits of alignment? Our analysis reveals that the preference collapse phenomenon (Xiao et al., 2024; Chakraborty et al., 2024), a known issue in alignment methods where models excessively favor certain responses, undesirably generalizes to multiple-choice questions, a crucial scenario where calibration performance can be conveniently evaluated. This generalization leads to overconfidence and poor calibration, as models tend to collapse their prediction to one option while ignoring others, regardless of the correctness of their chosen option. Building on these insights, we demonstrate that fine-tuning with domain-specific knowledge can extensively alleviate overconfidence on incorrect answers and marginally amplify overconfidence on correct answers, thus improving calibration performance.
For further analysis, we develop a theoretical framework that characterizes the conditions under which calibration can be restored. We introduce the concepts of calibratable and non-calibratable regimes, defined by upper and lower bounds of Expected Calibration Error (ECE) and a critical accuracy threshold. This framework provides insights into when and how models can achieve good calibration while maintaining performance. We find that preference-aligned models typically lie in the calibratable regime, and propose a calibration-aware fine-tuning (CFT) to restore calibration without compromising performance. However, when we overly fine-tune these models to achieve better performance, they shift into the non-calibratable regime. In this case, we identify a fundamental trade-off between ECE and performance, and develop an EM-algorithm-based ECE regularization for the CFT loss to effectively navigate this trade-off.
@@ 2 Related Work
Calibration for LLMs. Several studies have specifically examined LLM calibration (Jiang et al., 2021; Xiao et al., 2022; Chen et al., 2022), identifying miscalibration issues and evaluating various interventions. Some researchers have explored auxiliary models (Zhang et al., 2021; Kadavath et al., 2022; Shen et al., 2024) to improve calibration, while others have investigated prompt-based approaches for RLHF-tuned LLMs (Xiong et al., 2023). Other researchers teach the LLMs to express their uncertainty and calibration (Lin et al., 2022; Tian et al., 2023). A distinct line of research focuses on debiasing in-context learning (Abbas et al., 2024; Han et al., 2023; Jiang et al., 2023b), though this differs from statistical calibration. Notably, previous studies consistently identify temperature scaling as the most effective calibration method. Given its strong performance, we use temperature scaling as our primary baseline for comparison.
@@ 3 Preliminaries
To answer multiple-choice questions, LLMs typically process the question along with all available answer choices as a single input. The model generates a probability distribution over the possible responses, allowing selection of the most likely correct answer. To ensure that LLMs respond with a single letter from $\{A,B,C,D\}$, a hint–‘Answer with only a single letter’–is appended to the prompt. For models in the Llama family, we use an additional system prompt to enforce this response format. An example of this setup is illustrated in Table 1.
@@ 4.1 A Generative Model for Multiple-Choice Problems
Consider a multiple choice dataset be defined as $\mathcal{S}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$, which can be viewed as an examination created by a test designer. The designer’s main constraint is to ensure an approximately balanced distribution, with each option ($A,B,C,$ and $D$) containing about 25% of the correct answers. This test designer can be conceptualized as a probabilistic generative model that assigns correct answers to positions according to their implicit preferences while maintaining this distribution. For a given dataset, there exist many different probabilistic generative models. Here are two illustrative examples:
@@ 5 Algorithms
5.1 Overconfidence Issue of Aligned LLMs
In this session, we study why LLMs become poorly calibrated after preference alignment. We first discuss the preference collapse phenomenon (Xiao et al., 2024; Chakraborty et al., 2024) in alignment methods like RLHF and DPO. When LLMs are aligned with a sample $(x,y_{w},y_{l})$, they tend to collapse their relative preference on $y_{w}$ while ignoring $y_{l}$. This typically results in a preference ratio exceeding human preference proportions: $\pi(y_{w}|x)/(\pi(y_{w}|x)+\pi(y_{l}|x))>\mathbb{P}(y_{w}\succ y_{l}|x)$. This preference collapse issue generalizes to unseen preference pairs $(y_{1},y_{2})$, where LLMs collapse their preference to one option.
@@ 5.1 Overconfidence Issue of Aligned LLMs
5.2 EM-Algorithms framework for Probabilistic Model
@@ 6.1 Experimental Setup
Models.
We evaluate our methods using four open-source large language models: Llama-3.1-Tulu-8B (Lambert et al., 2024), Vicuna-7B-v1.5 (Chiang et al., 2023), Olmo 2-7B (OLMo et al., 2024), Mistral-7B (Jiang et al., 2023a), with detailed information provided in Appendix C.1.
@@
Baselines.
We validate the effectiveness of our methods by comparing performance across four methods:
@@
Table 2: Performance comparison among DPO/RLHF, Temperature Scaling, CFT, and RCFT across four models (Llama3.1-8B-Tulu, Vicuna-7B, Olmo2-7B, and Mistral-7B) in in-domain and out-domain scenarios. Best results in each metric block are bolded. Blue highlights indicate superior in-domain conf-ECE of our CFT while red highlights denote best in-domain accuracy of our RCFT. $\downarrow$”/$\uparrow$” means the smaller/larger the better. “-” means the results of Temperature Scaling are the same as the original DPO/RLHF version.
Model Method conf-ECE $\downarrow$ cw-ECE $\downarrow$ Accuracy $\uparrow$ Win Rate $\uparrow$
@@
Llama3.1- 8B-Tulu DPO 0.1953 0.1212 0.0953 0.0650 0.6228 0.7810 0.7295
Temp Scale. 0.1126 0.0679 0.0336 0.0514 - - -
CFT(Ours) 0.0239 0.0688 0.0582 0.0375 0.6410 0.8000 0.7460
RCFT(Ours) 0.0897 0.0810 0.0771 0.0526 0.8341 0.7991 0.7118
Vicuna-7B RLHF 0.1422 0.0852 0.0979 0.0560 0.4344 0.5233 0.2271
Temp Scale. 0.0598 0.0224 0.0488 0.0484 - - -
CFT(Ours) 0.0379 0.0331 0.0583 0.0491 0.4481 0.6172 0.2279
RCFT(Ours) 0.0474 0.0672 0.0459 0.0530 0.6015 0.6035 0.2257
Olmo2-7B DPO 0.1555 0.1325 0.0873 0.1331 0.6210 0.6635 0.7493
Temp Scale. 0.0665 0.1160 0.0355 0.1196 - - -
CFT(Ours) 0.0544 0.0225 0.0804 0.0637 0.6606 0.7085 0.7588
RCFT(Ours) 0.0989 0.0781 0.0806 0.0707 0.8510 0.7099 0.7517
Mistral-7B DPO 0.2010 0.1318 0.0909 0.1103 0.6331 0.7567 0.7066
Temp Scale. 0.0802 0.0991 0.0399 0.0909 - - -
CFT(Ours) 0.0651 0.0424 0.0712 0.0614 0.6514 0.7863 0.7124
RCFT(Ours) 0.0979 0.0731 0.0877 0.0739 0.8297 0.7768 0.7221
Evaluation Metrics.
In this paper, we use four metrics to evaluate the effectiveness of the proposed method. They are (1) class-wise Expected Calibration Error (cw-ECE), (2) Conference Expected Calibration Error (conf-ECE) (Guo et al., 2017), (3) Accuracy and (4) Win Rate where the first two are to evaluate the calibration performance (the lower the better). We use Accuracy and Win Rate to evaluate the ability of LLMs to understand languages and answer the questions correctly (the higher the better). The continuous cw-ECE and conf-ECE are defined in Def. 3.1 and 3.2. We use their discrete versions by stratifying all the samples into ten bins according to the prediction probability. The discrete versions can be found in Section C.1.
@@
Datasets.
We evaluate the ECE and Accuracy using four diverse datasets: MMLU, MedMCQA, OpenBookQA, Arc-Challenge, and evaluate the Win Rate using UltraFeedback-Binarized-Preferences. The detailed descriptions of these datasets are provided in Appendix C.1.
Evaluation Schemes.
We evaluate not only the in-domain performance (i.e., the training and testing data share the same domain), but also the out-domain (zero-shot) ability on an out-of-domain dataset. More specifically, we have two schemes:
@@ 6.2 Numerical Results
Table 2 compares four methods—DPO, Temperature Scaling, CFT (ours), and RCFT (ours)—across four language models (Llama3.1-8B-Tulu, Vicuna-7B, Olmo2-7B, and Mistral-7B) in terms of calibration performance (conf-ECE, cw-ECE) and language ability retention (Accuracy, Win Rate) under both in-domain and out-domain evaluations. Key insights include:
@@
CFT significantly improves calibration while preserving or enhancing language capabilities. For Vicuna-7B, CFT reduces in-domain conf-ECE by 73% compared to DPO (0.0379 vs. 0.1422) while increasing out-domain accuracy to 0.6172 (vs. DPO’s 0.5233). Similarly, for Mistral-7B, CFT achieves a 68% reduction in in-domain conf-ECE (0.0651 vs. DPO’s 0.201) and maintains strong out-domain accuracy (0.7863 vs. DPO’s 0.7567). CFT also outperforms Temperature Scaling in challenging scenarios, such as reducing Olmo2-7B’s out-domain cw-ECE to 0.0637 (vs. Temperature Scaling’s 0.1196), paired with improved accuracy (0.7085 vs. DPO’s 0.6635). This verifies that the DPO or RLHF models typically lie in the calibratable regime, their calibration can be restored without sacrificing the language capabilities.
@@
RCFT prioritizes accuracy gains while maintaining competitive calibration. For Llama3.1-8B-Tulu, RCFT achieves the highest in-domain accuracy (0.8341) compared to DPO (0.6228) but exhibits slightly higher calibration errors (in-domain conf-ECE: 0.0897 vs. CFT’s 0.0239). This trend holds for Olmo2-7B, where RCFT boosts in-domain accuracy to 0.851 (vs. DPO’s 0.621) with modest conf-ECE (0.0989). Here, the models move to the non-calibratable regime. While RCFT trades some calibration performance for accuracy, it remains comparable to Temperature Scaling, making it suitable for accuracy-critical applications.
Temperature Scaling reliably reduces calibration errors but fails to improve language abilities. For Vicuna-7B, it lowers in-domain conf-ECE to 0.0598 (vs. DPO’s 0.1422), yet accuracy and Win Rate remain unchanged (0.4344 and 0.5233, identical to DPO). This underscores its limitation as a post-hoc confidence-adjustment method that does not enhance model capability.
DPO (RLHF for Vicuna-7B), the baseline method, consistently underperforms in calibration (e.g., Mistral-7B’s in-domain conf-ECE: 0.201) and achieves moderate accuracy (e.g., Llama3.1-8B-Tulu’s out-domain accuracy: 0.781), highlighting the necessity of specialized calibration-focused techniques like CFT or RCFT.
Notably, CFT demonstrates robust out-domain generalization. For Llama3.1-8B-Tulu, it improves out-domain Win Rate to 0.746 (vs. DPO’s 0.7295) while maintaining low cw-ECE (0.0375). RCFT, though calibration-weaker, generalizes accuracy gains effectively, as seen in Mistral-7B’s out-domain Win Rate (0.7221 vs. DPO’s 0.7066).
In summary, CFT is optimal for calibration-sensitive tasks, whereas RCFT excels in accuracy-driven applications. Both methods outperform DPO/RLHF and Temperature Scaling, showcasing adaptability across diverse models and evaluation scenarios.

### v2

@@ 1 Introduction
In this paper, we conduct a systematic investigation into two fundamental questions: (1) Why does preference alignment affect calibration? and (2) How can we effectively restore calibration while maintaining the benefits of alignment? Our analysis reveals that the preference collapse phenomenon (Xiao et al., 2024a; Chakraborty et al., 2024), a known issue in alignment methods where models excessively favor certain responses, undesirably generalizes to multiple-choice questions, a crucial scenario where calibration performance can be conveniently evaluated. This generalization leads to overconfidence and poor calibration, as models tend to collapse their prediction to one option while ignoring others, regardless of the correctness of their chosen option. Building on these insights, we demonstrate that fine-tuning with domain-specific knowledge can extensively alleviate overconfidence on incorrect answers and marginally amplify overconfidence on correct answers, thus improving calibration performance.
For further analysis, we develop a theoretical framework that characterizes the conditions under which calibration can be restored. We introduce the concepts of calibratable and non-calibratable regimes, defined by upper and lower bounds of Expected Calibration Error (ECE) and a critical accuracy threshold. This framework provides insights into when and how models can achieve good calibration while maintaining performance. We consider fine-tuning the model after RLHF or DPO, rather than modifying these methods directly, as in many practical cases only the final DPO/RLHF model is available, while the pretrained or SFT model is not. We find that preference-aligned models typically lie in the calibratable regime, and propose a calibration-aware fine-tuning (CFT) approach that restores calibration without compromising performance. However, when we overly fine-tune these models to achieve better performance, they shift into the non-calibratable regime. In this case, we identify a fundamental trade-off between ECE and performance, and develop an EM-algorithm-based ECE regularization for the CFT loss to effectively navigate this trade-off.
@@ 2 Related Work
Calibration for LLMs. Several studies have specifically examined LLM calibration (Jiang et al., 2021; Xiao et al., 2022; Chen et al., 2022), identifying miscalibration issues and evaluating various interventions. Some researchers have explored auxiliary models (Zhang et al., 2021; Kadavath et al., 2022; Shen et al., 2024) to improve calibration, while others have investigated prompt-based approaches for RLHF-tuned LLMs (Xiong et al., 2023). Other researchers teach the LLMs to express their uncertainty and calibration (Lin et al., 2022; Tian et al., 2023; Xu et al., 2024b). A distinct line of research focuses on debiasing in-context learning (Abbas et al., 2024; Han et al., 2023; Jiang et al., 2023b), though this differs from statistical calibration. Researchers have studied improving the calibration of LLMs during the SFT phase (Han et al., 2024) as well as during the PPO/DPO phase (Tao et al., 2024; Leng et al., 2024). Notably, previous studies consistently identify temperature scaling as the most effective calibration method. Given its strong performance, we use temperature scaling as our primary baseline for comparison.
@@ 3 Preliminaries
To answer multiple-choice questions, LLMs typically process the question along with all available answer choices as a single input. The model generates a probability distribution over the possible responses, allowing selection of the most likely correct answer. To ensure that LLMs respond with a single letter from $\{A,B,C,D\}$ [^1] We use the four-option format $\{A,B,C,D\}$ for readability; however, the analysis extends to a general $k$-choice setting. , a hint—‘Answer with only a single letter’—is appended to the prompt. For models in the Llama family, we use an additional system prompt to enforce this response format. An example of this setup is illustrated in Table 1.
@@ 4.1 A Generative Model for Multiple-Choice Problems
Consider a multiple-choice dataset defined as $\mathcal{S}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$, which can be viewed as an examination created by a test designer. The designer’s main constraint is to ensure an approximately balanced distribution, with each option ($A,B,C,$ and $D$) containing about 25% of the correct answers. This test designer can be conceptualized as a probabilistic generative model that assigns correct answers to positions according to their implicit preferences while maintaining this distribution. For a given dataset, there exist many different probabilistic generative models. Here are two illustrative examples:
@@ 5 Algorithms
The objective in Equation (2) can be reformulated as minimizing a combination of accuracy loss and ECE loss using a Lagrange multiplier. In this section, we discuss how to approximate both losses specifically tailored for LLMs.
5.1 Approximating the Accuracy Loss for LLMs
In this session, we study why LLMs become poorly calibrated after preference alignment. We first discuss the preference collapse phenomenon (Xiao et al., 2024a; Chakraborty et al., 2024) in alignment methods like RLHF and DPO. When LLMs are aligned with a sample $(x,y_{w},y_{l})$, they tend to collapse their relative preference on $y_{w}$ while ignoring $y_{l}$. This typically results in a preference ratio exceeding human preference proportions: $\pi(y_{w}|x)/(\pi(y_{w}|x)+\pi(y_{l}|x))>\mathbb{P}(y_{w}\succ y_{l}|x)$. This preference collapse issue generalizes to unseen preference pairs $(y_{1},y_{2})$, where LLMs collapse their preference to one option.
@@ 5.1 Approximating the Accuracy Loss for LLMs
5.2 Approximating the ECE loss for LLMs
@@ 6.1 Experimental Setup
Models. We evaluate our methods using four open-source large language models: Llama-3.1-Tulu-8B (Lambert et al., 2024), Vicuna-7B-v1.5 (Chiang et al., 2023), Olmo 2-7B (OLMo et al., 2024), Mistral-7B (Jiang et al., 2023a), with detailed information provided in Appendix C.1.
@@
Baselines. We validate the effectiveness of our methods by comparing performance across five methods:
@@
  • Label Smoothing (Müller et al., 2019): Replaces hard one-hot labels with a smooth label that assigns $1-\varepsilon$ to the correct class and distributes $\varepsilon$ over the remaining classes.
@@
Table 2: Performance comparison among DPO/RLHF, Temperature Scaling, Label Smoothing, CFT, and RCFT across four models (Llama3.1-8B-Tulu, Vicuna-7B, Olmo2-7B, and Mistral-7B) in in-domain and out-domain scenarios. Best results in each metric block are bold. Blue highlights indicate superior in-domain conf-ECE of our CFT while red highlights denote best in-domain accuracy of our RCFT. $\downarrow$”/$\uparrow$” means the smaller/larger the better. “-” means the results of Temp. Scale. are the same as the original DPO/RLHF version.
Model Method conf-ECE $\downarrow$ cw-ECE $\downarrow$ Accuracy $\uparrow$
@@
Llama3.1- 8B-Tulu DPO 0.1953 0.1212 0.0953 0.0650 0.6228 0.7810
Temp. Scale. 0.1126 0.0679 0.0336 0.0514 - -
Label Smooth. 0.1898 0.1009 0.0692 0.0639 0.6372 0.7116
CFT(Ours) 0.0239 0.0688 0.0582 0.0375 0.6410 0.8000
RCFT(Ours) 0.0897 0.0810 0.0771 0.0526 0.8341 0.7991
Vicuna-7B RLHF 0.1422 0.0852 0.0979 0.0560 0.4344 0.5233
Temp. Scale. 0.0598 0.0224 0.0488 0.0484 - -
Label Smooth. 0.1221 0.0823 0.0517 0.0544 0.4517 0.5767
CFT(Ours) 0.0379 0.0331 0.0583 0.0491 0.4481 0.6172
RCFT(Ours) 0.0474 0.0672 0.0459 0.0530 0.6015 0.6035
Olmo2-7B DPO 0.1555 0.1325 0.0873 0.1331 0.6210 0.6635
Temp. Scale. 0.0665 0.1160 0.0355 0.1196 - -
Label Smooth. 0.1010 0.0499 0.0791 0.1298 0.6808 0.6431
CFT(Ours) 0.0544 0.0225 0.0804 0.0637 0.6606 0.7085
RCFT(Ours) 0.0989 0.0781 0.0806 0.0707 0.8510 0.7099
Mistral-7B DPO 0.2010 0.1318 0.0909 0.1103 0.6331 0.7567
Temp. Scale. 0.0802 0.0991 0.0399 0.0909 - -
Label Smooth. 0.1874 0.1121 0.0900 0.0990 0.6479 0.6997
CFT(Ours) 0.0651 0.0424 0.0712 0.0614 0.6514 0.7863
RCFT(Ours) 0.0979 0.0731 0.0877 0.0739 0.8297 0.7768
Evaluation Metrics. In this paper, we use four metrics to evaluate the effectiveness of the proposed method. They are (1) class-wise Expected Calibration Error (cw-ECE), (2) Conference Expected Calibration Error (conf-ECE) (Guo et al., 2017), (3) Accuracy and (4) Win Rate where the first two are to evaluate the calibration performance (the lower the better). We use Accuracy and Win Rate to evaluate the ability of LLMs to understand languages and answer the questions correctly (the higher the better). The continuous cw-ECE and conf-ECE are defined in Def. 3.1 and 3.2. We use their discrete versions by stratifying all the samples into ten bins according to the prediction probability. The discrete versions can be found in Section C.1.
@@
Datasets. We evaluate the ECE and Accuracy using four diverse datasets: MMLU, MedMCQA, OpenBookQA, Arc-Challenge, and evaluate the Win Rate using AlpacaEval, Arena-Hard, and UltraFeedback-Binarized-Preferences. The detailed descriptions of these datasets are provided in Appendix C.1.
Evaluation Schemes. We evaluate not only the in-domain performance (i.e., the training and testing data share the same domain), but also the out-domain (zero-shot) ability on an out-of-domain dataset. More specifically, we have two schemes:
@@
Table 3: Win rate comparisons among DPO/RLHF (DPO used in Table 3), CFT and RCFT across four models (Llama3.1-8B-Tulu, Vicuna-7B, Olmo2-7B and Mistral-7B) on three datasets (AlpacaEval, Arena-Hard and Ultrafeedback). The best performance for each dataset is in bold. The competitive performance indicates that our methods can preserve the alignment performance.
Model AlpacaEval (vs DPO) AlpacaEval Arena-Hard Ultrafeedback
CFT vs DPO RCFT vs DPO DPO CFT RCFT DPO CFT RCFT DPO CFT RCFT
Llama-3.1-8B-Tulu 51.68 vs 48.32 46.83 vs 53.16 21.4 22.6 19.6 44.6 45.0 43.6 0.7295 0.7460 0.7118
Vicuna-7B 46.46 vs 53.54 50.43 vs 49.57 2.60 2.60 3.60 1.00 1.00 1.00 0.2271 0.2279 0.2257
Olmo2-7B 62.48 vs 37.52 46.12 vs 53.88 24.2 22.9 23.1 19.4 19.2 20.2 0.7493 0.7588 0.7517
Mistral-7B 46.96 vs 53.04 49.81 vs 50.19 26.0 26.8 25.2 18.9 18.3 18.0 0.7066 0.7124 0.7221
@@ 6.2 Numerical Results
Table 2 compares five alignment methods—DPO/RLHF, Temperature Scaling, Label Smoothing, CFT (ours), and RCFT (ours)—across four language models (Llama3.1-8B-Tulu, Vicuna-7B, Olmo2-7B, and Mistral-7B) in terms of calibration performance (conf-ECE, cw-ECE) and Accuracy under both in-domain and out-domain evaluations. [^2] Code is publicly available at https://github.com/ZhanliangAaronWang/RestoreLLMCalibration. Key insights include:
@@
CFT significantly improves calibration while preserving or enhancing language capabilities. For Vicuna-7B, CFT reduces in-domain conf-ECE by 73% compared to DPO (0.0379 vs. 0.1422) while increasing out-domain accuracy to 0.6172 (vs. DPO’s 0.5233). Similarly, for Mistral-7B, CFT achieves a 68% reduction in in-domain conf-ECE (0.0651 vs. DPO’s 0.201) and maintains strong out-domain accuracy (0.7863 vs. DPO’s 0.7567). CFT also outperforms Temperature Scaling and Label Smoothing in challenging scenarios, such as reducing Olmo2-7B’s out-domain conf-ECE to 0.0225 (vs. Label Smoothing’s 0.0499 and Temperature Scaling’s 0.1160). This verifies that DPO or RLHF-aligned models typically lie in the calibratable regime, where calibration can be restored without sacrificing language capabilities.
@@
RCFT prioritizes accuracy gains while maintaining competitive calibration. For Llama3.1-8B-Tulu, RCFT achieves the highest in-domain accuracy (0.8341) compared to DPO (0.6228) but exhibits slightly higher calibration errors (in-domain conf-ECE: 0.0897 vs. CFT’s 0.0239). This trend holds for Olmo2-7B, where RCFT boosts in-domain accuracy to 0.851 (vs. DPO’s 0.621 and Label Smoothing’s 0.6808) with modest conf-ECE (0.0989). Here, the models move to the non-calibratable regime. While RCFT trades some calibration performance for accuracy, it remains comparable to Temperature Scaling, making it suitable for accuracy-critical applications.
Baseline methods show consistent limitations. Temperature Scaling reduces calibration errors (e.g., Vicuna-7B conf-ECE: 0.0598 vs. 0.1422) but cannot improve accuracy (marked by "-"). Label Smoothing provides moderate calibration benefits but underperforms our methods—for example, on Mistral-7B, its in-domain conf-ECE (0.1874) is significantly higher than CFT’s (0.0651), and its accuracy gains (+1.5% over DPO) are dwarfed by RCFT’s +31%. DPO/RLHF consistently shows the poorest calibration (e.g., Mistral-7B conf-ECE: 0.201), highlighting the need for specialized techniques.
CFT and RCFT preserve alignment quality across diverse benchmarks. Table 3 demonstrates this through three complementary evaluation protocols:
1. AlpacaEval Head-to-Head (CFT/RCFT vs DPO): Responses from our methods and the original DPO model were directly compared by GPT-4 (judge):
  • Protocol: Same prompts → CFT/RCFT vs DPO responses → GPT-4 preference
  • Results: Statistical parity (e.g., Vicuna-7B: 50.43% RCFT vs 49.57% DPO)
2. Standard AlpacaEval (vs GPT-4o): Methods were evaluated against the standard GPT-4o baseline:
  • Protocol: Responses vs GPT-4o → GPT-4 preference
  • Results: CFT achieved best performance for Llama3.1 (22.6%) and Mistral-7B (26.8%)
3. Arena-Hard (vs GPT-4): Responses were compared against the standard GPT-4 baseline [^3] The value of the instruct version of Llama-3.1-8B is a 20.6%. It is not clear why the Tulu version has a significant improvement.:
  • Protocol: Responses vs GPT-4 → GPT-4o preference
  • Results: RCFT achieved best for Olmo2-7B (20.2%); CFT showed gains (Llama3.1: 45.0% vs DPO’s 44.6%)
4. Ultrafeedback Binary Selection: Models performed preference classification:
  • Protocol: Given question, response A, response B → Select better response
  • Results: CFT improved win rates for all models (e.g., Olmo2-7B: 0.7588 vs DPO’s 0.7493)
In summary, CFT excels in calibration-sensitive tasks while RCFT dominates accuracy-critical applications. Both methods outperform existing techniques (including Label Smoothing), preserve alignment quality across all benchmarks, and demonstrate robust generalization across models and evaluation scenarios.